### v1

# Measurement of zero-frequency fluctuations generated by coupling between Alfvén modes in the JET tokamak

Abstract
We report the first experimental detection of a zero-frequency fluctuation that is pumped by an Alfvén mode in a magnetically confined plasma. Core-localized bidirectional Alfvén modes of frequency inside the toroidicity-induced gap (and its harmonics) exhibit three-wave coupling interactions with a zero-frequency fluctuation. The observation of the zero-frequency fluctuation is consistent with theoretical and numerical predictions of zonal modes pumped by Alfvén modes, and is correlated with an increase in the deep core ion temperature, temperature gradient, and confinement factor $H_{89,P}$. Despite the energetic particle transport induced by the Alfvén eigenmodes, the generation of a zero-frequency fluctuation that can suppress the turbulence leads to an overall improvement of confinement.
preprint: APS/123-QED
Achieving net energy gain from magnetically confined thermonuclear fusion is primarily hindered by the turbulent transport of heat and particles across the confining magnetic field. Turbulent transport determines the background plasma profiles, and, as a result, the overall fusion power-balance in a reactor. In a burning plasma, fusion reactions between deuterium and tritium (DT) will generate supra-thermal alpha particles of energy $E=3.5$ MeV, more than two orders of magnitude higher than the temperature of the thermal plasma, typically $T\approx 10-15$ keV. Fusion-born alpha particles are expected to have characteristic velocities larger than the Alfvén speed $v_{A}=B/(\mu_{0}\sum_{i}m_{i}n_{i})^{1/2}$, where $B$ is the background magnetic field, $\mu_{0}$ is the vacuum magnetic permeability and the sum is taken over the thermal ion species of mass $m_{i}$ and density $n_{i}$. Such a super-Alfvénic population is capable of triggering magnetohydrodynamic (MHD) Alfvén instabilities, which have been predicted [1] and observed in a wide range of magnetic-confinement devices [2, 3, 4, 5, 6, 7]. For many years, the presence of fast-ion-driven Alfvén instabilities has been thought to be detrimental to the overall plasma confinement, since they are detrimental for the confinement of the energetic particles themselves [8, 9, 10]. Recently, however, numerical simulations of the turbulence have shown that fast-ion-driven Alfvénic activity can have a positive effect on the confinement of the thermal plasma [11, 12, 13, 14, 15]. The current leading hypothesis explaining this effect is the pumping by the Alfvén eigenmode of a zonal-flow component, which could interact with, and suppress, the turbulence. This seems to be in agreement with analytical predictions of zonal-flow excitation by Alfvén eigenmodes [16, 17, 18, 19] and with numerical simulations [20, 21, 22, 23, 24]. Zonal modes are zero-frequency perturbations that are constant within a magnetic-flux surface. In the traditional zonal-flow paradigm, a zonal flow generated by the turbulence fluctuations can produce stabilization [25, 26, 27, 28, 29]. The zonal-flow generation by Alfvén eigenmodes, however, has so far lacked direct experimental confirmation. In this article, we report the first experimental detection of a zero-frequency fluctuation that is pumped by Alfvén eigenmodes.
The Joint European Torus (JET [30]) is the magnetic-confinement facility that has come the closest to approaching the physics regimes of a burning plasma [31, 32, 33, 34]. The JET discharge 97090 was an L-mode characterized by a plasma current of $I_{p}=2.4$ MA, toroidal magnetic field $B_{0}=3.2$ T, line-integrated density $\int n_{e}\text{d}l=7.5\times 10^{19}$ m-2, Greenwald fraction of $f_{GW}=0.36$ [35] and $q_{95}=3.75$, all of which remained constant within uncertainties throughout the discharge. The central electron density was $n_{e}\approx 4\times 10^{19}$ m-3. The plasma was ion-cyclotron-resonance heated (ICRH) using the three-ion scheme [36, 37, 38] (main ions were composed of $80\%$ hydrogen and $20\%$ deuterium). A first Ohmic phase was followed by phases of increasing ICRH power of $P_{\text{ICRH}}=$ 2 MW (low), 4 MW (medium) and 7 MW (high, see black curve in figure 1.(a). A trace population of 3He ions absorbed $\gtrsim 90\%$ of the ICRH power and was accelerated to energies of 4-5 MeV, as predicted by TRANSP [39] and TORIC [40] (consistent with $\gamma$-ray measurements in similar discharges [41]). In the absence of tritium, MeV-range 3He can mimic the effect of fusion-born alpha particles in a burning plasma (although the anisotropic ICRH fast-ion distribution function is qualitatively different from the isotropic alpha-particle fusion source, and can result in differences in the Alfvénic-instability drive [42]). The fast 3He ions deposited $\approx 90\%$ of their energy on the thermal electrons by collisionally slowing down. The remainder of the ICRH power is predicted to be absorbed directly by electrons, with less than $1\%$ of the power predicted to be absorbed directly by the thermal ions. Thus, $\gtrsim 90\%$ of the heating power was predicted to be ultimately deposited on the electrons. As we will see, this is at odds with the increase in the deep-core ion temperature observed in this plasma.
Figure 1: (a) Magnetic spectrogram from the MHD magnetic-coil diagnostic in JET, along with the total ICRH power $P_{\text{ICRH}}$. (b) Electron temperature from the ECE diagnostic on axis (black) and at $\rho\approx 0.3$ (red). Also shown are $\tau_{ei}/\tau_{E}$ (cyan), the electron-ion collision time divided by the energy confinement time (TRANSP), and, the L-mode confinement factor $H_{89,P}$ (blue). Vertical dashed lines indicate times for which DBS measurements are shown in figures 3, 4 and 5.
The plasma exhibits internal sawtooth oscillations [43] with an inversion radius at $\rho\approx 0.3$ (consistent with the safety factor $q=1$), where $\rho$ is the magnetic-flux-surface label defined at the square root of the normalized toroidal magnetic flux ($\rho=0$ corresponds to the magnetic axis, while $\rho=1$ is the location of the edge separatrix). The sawtooth period increases with $P_{\text{ICRH}}$, as shown by the electron temperature $T_{e}$ in figure 1.(b) (from the electron-cyclotron-emission diagnostic, ECE), and consistent with the stabilization of sawteeth by fast ions [44]. This gives rise to the so-called monster sawteeth during the medium- and high-$P_{\text{ICRH}}$ phases. Following a sawtooth event, the on-axis electron temperature of $T_{e}\approx 5$ keV is shown to drop down to $\approx 2$ keV. The impact of the sawtooth crash is very weakly present near the inversion radius. In this manuscript, we analyze the plasma during time windows when the profiles have fully recovered after a sawtooth crash (dashed lines in figure 1.(b).
After recovery of the profiles from a sawtooth crash, the trapped 3He population at MeV energies was observed to destabilize a range of Alfvén eigenmodes during the medium- and high-$P_{\text{ICRH}}$ phases. Both phenomena are often simultaneously observed in tokamak experiments [45, 46, 47]. Figure 1.(a) shows the magnetic spectrogram of fluctuations from the MHD magnetic-coil diagnostic. During the medium-$P_{\text{ICRH}}$ phase, magnetic fluctuations associated with Alfvén modes are observed around $300$ kHz. During the high-$P_{\text{ICRH}}$ phase, separate branches are observed to sweep down in frequency in the range $300\rightarrow 280$ kHz, consistent with the expected temporal evolution of the central safety factor following a sawtooth crash [48, 49]. This behavior has already been observed in many major tokamaks, e.g., JT-60U [50, 48, 51, 49, 52], TFTR [53], and JET [54, 55, 47]. The modes are spatially localized around the $q=1$ surface with very low magnetic shear, as suggested by Doppler-backscattering measurements, pressure-constrained EFIT (including fast-ion pressure), and ECE.
The unstable Alfvén eigenmodes in figure 1.(a) have toroidal mode numbers in the range $n=3-6$. The toroidicity-induced gap frequency associated with the toroidal Alfvén eigenmode (TAE [56]) for large-aspect-ratio tokamaks is $f_{A}\approx v_{A}/4\pi qR_{0}$, where $R_{0}$ the major radius of the flux-surface centre. The numerical values for the density, $q=1$ and $R_{0}\approx 2.9$ m lead to $f_{A}\approx 270$ kHz. This is consistent with the magnetic fluctuations seen in figure 1.(a), which suggests that these modes live in the TAE gap, and could be identified as tornado TAEs [49]. Note that the plasma beta in this discharge is low ($\beta_{e}\lesssim 0.004$ for electrons), which is consistent with the conditions to excite shear Alfvén modes such as TAEs.
Figure 2: (a) Profiles of the main-ion temperature $T_{D}$ at low, medium, and high $P_{\text{ICRH}}$. (b) Normalized temperature-gradient scale length $a/L_{TD}$. Both $T_{D}$ and $a/L_{TD}$ increase inside $\rho\approx 0.4$ with increasing $P_{\text{ICRH}}$, despite negligible main-ion heating.
The dominant electron heating in this discharge increases the electron temperature at the inner radii from 2 keV (Ohmic) to 5 keV (high $P_{\text{ICRH}}$). The main-ion temperature in the deep core also increases (figure 2.(a) and approaches the electron temperature at medium and high $P_{\text{ICRH}}$. The measurements are made with main-ion charge-exchange recombination spectroscopy [57] during $10$ ms neutral-beam-injection (NBI) blips (performed every second), which allow the temperature to be measured without affecting the overall plasma. The increase in the main-ion temperature shows that it is not clamped at a specific value, as in recent experiments of pure electron-cyclotron-resonance heating [58, 59]. This is a very positive result for fusion, since burning plasmas will have dominant electron heating from the slowing down of 3.5 MeV alpha particles, but will require high main-ion temperatures for fusion reactions to occur. The main ions here also exhibit an increase in the normalized temperature gradient $a/L_{TD}=-\text{d}\log T_{D}/\text{d}\rho$ inside $\rho\approx 0.4$ ($a$ is the plasma minor radius). The increase in $a/L_{TD}$ starts outside the sawtooth inversion radius at $\rho\approx 0.3$, and suggests that fast-ion-driven modes may be improving the energy confinement of the main ions.
The increase in the main-ion temperature is a striking finding. The main ions experience negligible direct heating from ICRH, and their collisional coupling to the electrons decreases as the temperature increases. This is demonstrated in figure 1.(b), where the electron-ion collision time $\tau_{ei}$ divided by the total energy-confinement time $\tau_{E}$ is shown to increase (as modeled by TRANSP). Linear Landau damping of the TAE on the main ions is also expected to be very weak due to low ion beta [60, 61], as is nonlinear Landau damping [62, 63, 64, 65] (here $\beta_{i}\lesssim 0.003$). This suggests that a different mechanism must explain the ion-temperature increase. The total energy-confinement factor for the L-mode is shown to be $H_{89,P}<1$ in the low-$P_{\text{ICRH}}$ case, but increases to $H_{89,P}>1$ with increasing $P_{\text{ICRH}}$ (figure 1.(b)). We conjecture that the improved confinement and ion-temperature increase observed here are due to the suppression of turbulent transport by zero-frequency zonal modes pumped by Alfvén eigenmodes [66, 12, 11, 14]. Below, we provide direct experimental evidence for the presence of such a mechanism. We stress that this situation is very different from NBI-heated discharges that use lower-energy fast ions, where most of the heating is absorbed by the thermal ions, external torque drives flow-shear suppression of the turbulence, and Alfvén modes are more difficult to destabilize. Similar experiments in JET with MeV-range ions have exhibited an improvement in the confinement of the thermal plasma [41, 67, 12].
To confirm that the Alfvén modes shown in figure 1 play a crucical role in the overall confinement, we performed turbulence measurements at $\rho\approx 0.25-0.35$ using the Doppler-backscattering system (DBS) in JET, which detects fluctuations of the electron density. In the Doppler-backscattering technique [68, 69, 70, 71, 72, 73, 74], a beam of microwaves is launched into the core plasma until it encounters a cutoff surface. The detector receives backscattered radiation from turbulent fluctuations with characteristic wavenumber $k_{\perp}$ that are local to the vicinity of the cutoff [75], as predicted by beam-tracing numerical simulations of the microwave beam [76]. We normalize the scattered turbulence wavenumber $k_{\perp}$ by the local ion sound gyroradius $\rho_{s}=c_{s}/\Omega_{D}$, where $c_{s}=\sqrt{T_{e}/m_{D}}$ is the sound speed, $m_{D}$ is the deuterium mass, $\Omega_{D}=q_{D}B_{0}/m_{D}c$ its gyro-frequency, $q_{D}$ its charge, and $c$ the speed of light. The DBS measurements presented here are sensitive to density fluctuations at wavenumbers $k_{\perp}\rho_{s}\approx 2-4$. Dedicated analysis of the scattering trajectory using the beam-tracing code Scotty [76, 75] shows that $k_{\perp}$ has a small toroidal component that overlaps with $n=3-6$ of the unstable Alfvén modes in figure 1.(a). This suggests that DBS measurements can be sensitive to Alfvén modes in these conditions.
Figures 3, 4 and 5 show the power spectrum $P_{s}(f)$ and the bicoherence $B_{s}(f_{1},f_{2})$ of the electron-density fluctuations for the low-, medium- and high-$P_{\text{ICRH}}$ phases, respectively. The measurements are taken at specific times that are shown by the colored vertical dashed lines in figure 1.(b). The backscattered power is calculated from the Fourier-transformed backscattered complex signal amplitude $\hat{A}_{j}$ as $P_{s}(f)=\langle|\hat{A}_{j}|^{2}\rangle_{T}$, where the .$\langle.\rangle_{T}$ denotes an ensemble average and the $f=0$ component of $\hat{A}_{j}$ is removed ($j$ denotes the sample realization). The period of each time window is $T=0.5-1$ ms, which is much longer than the turbulence correlation time, which scales as $a/c_{s}\sim 10^{-5}-10^{-6}$ s. The bicoherence [77] associated with the ensemble of realization signals $\{A_{j}\}$ is
$B_{s}(f_{1},f_{2})=\frac{\langle\hat{A}_{j}(f_{1})\hat{A}_{j}(f_{2})\hat{A}_{j}(f_{1}+f_{2})^{*}\rangle_{T}}{\langle|\hat{A}_{j}(f_{1})\hat{A}_{j}(f_{2})|^{2}\rangle_{T}^{\frac{1}{2}}\langle|\hat{A}_{j}(f_{1}+f_{2})|^{2}\rangle_{T}^{\frac{1}{2}}}.$ (1)
The modulus of the bicoherence $|B_{s}(f_{1},f_{2})|$ in equation (1) is high when the frequency $f_{1}$ has a phase-matched relationship with the frequency $f_{2}$ to generate a third frequency at $f_{1}\pm f_{2}$. The ensemble averaging .$\langle.\rangle_{T}$ is critical to eliminate independent modes at the frequencies $f_{1}\pm f_{2}$. The bicoherence is a measure of three-wave nonlinear coupling between fluctuations of frequency $f_{1}$ and $f_{2}$.
Figure 3 shows the modulus of the bicoherence $|B_{s}(f_{1},f_{2})|$ and the power spectrum $P_{s}(f)$ at low $P_{\text{ICRH}}$. The bicoherence is calculated for $255$ windows of period $T=1$ ms that are half-overlapping. The DBS beam reaches the cutoff surface at $\rho=0.247$ (where $q\approx 1$), corresponding to the scattered wavenumber $k_{\perp}\rho_{s}=2.58$. The power spectrum $P_{s}(f)$ is broadband for fluctuations in the range $f\lesssim 200$ kHz characteristic of ion-temperature-gradient-driven turbulence (ITG). A peak at zero frequency is observed, consistent with many previous DBS measurements. The bicoherence displays broadband phase-matched relationships between frequencies in the range $f_{1},f_{2}<100$ kHz, as expected from nonlinear interactions between turbulent eddies characteristic of drift-wave turbulence. The amplitude, however, remains low at $\approx 0.2$, and this suggests that nonlinear interactions between low-frequency drift-wave fluctuations are broadband and modest in the low-$P_{\text{ICRH}}$ phase.
Figure 3: (Left) Bicoherence spectrum at low $P_{\text{ICRH}}$. (Right) Power spectrum $P_{s}(f)$. Both are characteristic of broadband drift-wave turbulence.
Figure 4 shows the modulus of the bicoherence and the power spectrum at medium $P_{\text{ICRH}}$, calculated for a total of $63$ half-overlapping windows of period $T=0.5$ ms. The DBS beam reaches the cutoff surface at $\rho=0.355$ ($q\approx 1$), corresponding to the scattered wavenumber $k_{\perp}\rho_{s}=2.07$. The power spectrum $P_{s}(f)$ is broader than in the low-$P_{\text{ICRH}}$ phase, featuring fluctuations in the range $f\lesssim 800$ kHz, as well as spectral peaks at the local Alfvén gap frequency $f_{A}\approx 290$ kHz and a harmonic at $2f_{A}$. The spectral peaks appear for positive and negative frequencies, suggesting the presence of bidirectional Alfvénic activity. Careful analysis of the DBS microwave beam, pressure-constrained EFIT, and the $q=1$ constraint from ECE shows that the modes observed in figure 4 are well localized around $q=1$. This is consistent with previous observations of core-localized TAEs in TFTR [78, 79, 80] and tornado modes in JT-60U [49]. Such modes were predicted to exist at low shear [81], consistent with the low shear value in this JET experiment ($\hat{s}\lesssim 0.1$). The bicoherence in figure 4 displays broadband phase-matched relationships for multiples of the local Alfvén gap frequency $f_{A}$, the strongest of which are $(f_{1},f_{2})=\{(\pm f_{A},\pm f_{A})\}$ (amplitude $\approx 0.6$), which correspond to co-propagating TAEs that couple to generate a perturbation of frequency ±$\pm 2f_{A}$, as well as $(f_{1},f_{2})=\{(2f_{A},-f_{A}),(-f_{A},2f_{A})\}$, which correspond to perturbations of frequency $2f_{A}$ coupling to a counter-propagating perturbation at frequency $f_{A}$ that generates a co-propagating perturbation at $f_{A}$. The bicoherence also clearly exhibits phase-matched relationships between perturbations at multiples of $f_{A}$ and a zero-frequency fluctuation. To our knowledge, this is the first experimental demonstration in a magnetically confined plasma that Alfvén modes living inside the fundamental toroidicity-induced frequency gap (and higher-frequency gaps) beat with counter-propagating perturbations to generate a zero-frequency fluctuation. The zero-frequency fluctuations observed here can naturally be interpreted as zonal modes driven by Alfvén eigenmodes, which have been predicted analytically [16, 17, 18, 19] and numerically [20, 22, 23, 24, 12, 11, 14]. Such zonal modes can suppress ITG-driven turbulence, which would explain the increase of the ion temperature and its gradient (figure 2).
Figure 4: The same as figure 3, but at medium $P_{\text{ICRH}}$. (Left) The bicoherence displays phase-matched relationships for multiples of $f_{A}$ (vertical dashed lines). The strongest interactions are between frequencies ±$\pm f_{A}$ and ±$\pm 2f_{A}$. A weaker phase-matched relationship is observed between fluctuations at multiples of $f_{A}$ and a zero-frequency fluctuation. (Right) The power spectrum $P_{s}(f)$ displays spectral peaks at ±$\pm f_{A}$ and ±$\pm 2f_{A}$.
Figure 5 shows the modulus of the bicoherence and the power spectrum at high $P_{\text{ICRH}}$, calculated for a total of $127$ half-overlapping windows of period $T=1$ ms. The DBS beam reaches the cutoff surface at $\rho=0.251$ ($q\approx 0.92$) for $k_{\perp}\rho_{s}=3.78$. The power spectrum is broadband for fluctuations in the range $f\lesssim 600$ kHz, but also displays spectral peaks at the local Alfvén gap frequency $f_{A}\approx 290$ kHz and its multiples. Similarly to figure 4, the bicoherence displays broadband phase-matched relationships for multiples of $f_{A}$. Importantly, the highest-amplitude phase matching is now for the frequency pairs $(f_{1},f_{2})=\{(2f_{A},0),(2f_{A},-2f_{A}),(3f_{A},0),(3f_{A},-3f_{A})\}$ (amplitude $\approx 0.5$), i.e., the pairs that involve a zero-frequency fluctuation. This is different from the medium-$P_{\text{ICRH}}$ phase, in which the phase-matched relationships involving the zero-frequency fluctuation were subdominant to other mode-mode interactions. Interestingly, the modes at the fundamental Alfvén gap frequency $f_{A}$ display a weaker phase-matched relationship with the zero-frequency fluctuation. Higher multiples of the Alfvén frequency show weaker phase-matched relationships, resulting in a peculiar grid structure.
Figure 5: (Left) The bicoherence spectrum at high-$P_{\text{ICRH}}$ displays phase-matched relationships for multiples of the local Alfvén gap frequency $f_{A}$. The strongest phase matching is for frequency pairs involving a zero-frequency fluctuation: $(f_{1},f_{2})=\{(\pm 2f_{A},0),(\pm 3f_{A},0)\}$. (Centre) Zoom into the left pannel. (Right) The power spectrum $P_{s}(f)$ displays spectral peaks at $f_{A}$ and its multiples.
Figures 4 and 5 confirm that Alfvén modes living inside the toroidicity-induced Alfvén-frequency gap (and higher-frequency gaps) beat with counter-propagating Alfvén modes to generate a strong zero-frequency fluctuation that is consistent with a zonal mode. When this behavior is observed, the total energy-confinement factor increases ($H_{89,P}>1$), as do the main-ion temperature and its gradient in the deep core, despite negligible direct ion heating. This shows the strong beneficial effect that Alfvén eigenmodes can have on the confinement of the thermal plasma. This is to be constrasted with the stable-Alfvén-mode case (low $P_{\text{ICRH}}$), which exhibited low ion temperature and temperature gradient, and $H_{89,P}<1$. This is the first experimental confirmation in a magnetically confined plasma of a zero-frequency fluctuation that is pumped by Alfvén modes, and is correlated with hotter-than-expected main ions. It suggests that zonal modes driven by Alfvén modes can suppress the turbulence and lead to an overall improvement of energy confinement in the plasma.
Admittedly, the presence of unstable Alfvén eigenmodes is not always correlated with improved confinement. It is also responsible for strong energetic-particle transport, which is not studied here, and is highly detrimental to energy confinement. The balance between this and the improved confinement for the thermal plasma observed here is still an open problem. It seems plausible that if this balance is quantitatively understood, unstable Alfvén modes could be tailored to improve the overall confinement in the burning plasmas expected to be achieved in the imminent future.
This work has been supported by the Engineering and Physical Sciences Research Council (EPSRC) [EP/W026341/1]. The work of M.B. and A.A.S. was supported in part by the EPSRC grant EP/R034737/1. The work of A.A.S. was also supported in part by the STFC grant STW000903/1 and by the Simons Foundation via a Simons Investigator Award. V. H. Hall-Chen was partly funded by an A*STAR Green Seed Fund, C231718014, and a YIRG, M23M7c0127. This work was supported by the U.S. Department of Energy under contract number DE-AC02-09CH11466. The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.